OpenStudy (jessiegonzales):

Medal!! Please help!!! 1. Find the circumcenter of the triagle.

OpenStudy (jessiegonzales):

My answer is D! (-4,-2)

OpenStudy (jessiegonzales):

@jim_thompson5910 @IMStuck

OpenStudy (tkhunny):

Well, it has to pass through those three vertices. Have you considered the perpendicular bisectors of the three sides. The vertical and horizontal would be easiest.

OpenStudy (jessiegonzales):

yes :) I was making sure my answer was correct A. (-5/2,-4) B. (-5/2,-2) C. (-2,-5/2) D. (-4,-2) <---

OpenStudy (tkhunny):

What answer did you get and how did you get it?

OpenStudy (tkhunny):

What's half way across that base? It's not -4. That's one of the endpoints.

OpenStudy (jessiegonzales):

-2?

OpenStudy (tkhunny):

Don't guess. Prove it! \(\dfrac{(-1)+(-4)}{2} = -\dfrac{5}{2}\)

OpenStudy (jessiegonzales):

sorry okay? next step

OpenStudy (tkhunny):

What next step? You're done. (-5/2,-2)